OpenStudy (anonymous):

Differential Equations Problem. Can someone help me just through the steps of solving this DE x^2y''-3xy'+4y=x^2ln(x) I need the general solution

OpenStudy (anonymous):

First I found the Yc (complementary solution using Caushy Euler Equation and i got yc=C1.e^2x + C2.xe^2x and i just dont know what to do next (usually i solve it using the variation of parameter) but this time it doesnt seem to work! i think i need to substitute x=e^t but i am unable to figure out where to and when.

OpenStudy (anonymous):

ok I will do this on paper and then attach it, one second.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

\[ A x^2+2 B x^2 \ln (x)+\frac{1}{6} x^2 \ln ^3(x) \]

OpenStudy (anonymous):

and what is that?

OpenStudy (anonymous):

That is the general solution. You first try to find a general solution for x^2y''-3xy'+4y=0 in the form y=x^r you find that r= 2 is a double root. and you continue from there.

OpenStudy (anonymous):

yes that what i did (i said that i found the complementary solution using Cauchy Euler Equation ) which is exactly like this! but what did you do after that?

OpenStudy (anonymous):

the cauchy euler solution that you got is wrong.. just take a look at the solution i did, you also have to use variation of parameters. ( in my solution the homogenous solution is the comlemetary solution). Good luck.

OpenStudy (anonymous):

what I just did is the correct method, I am not sure if I did any calculation mistakes.. so you should go over my solution and redo it.

OpenStudy (anonymous):

Thank you very much! but it is the same cauchy euler solution double root =2 again thank you

OpenStudy (anonymous):

ohh you are right it is wrong! i left out the ln

OpenStudy (anonymous):

ok no problem!

OpenStudy (anonymous):

Sir you are a boss!

OpenStudy (anonymous):

After finding that x^2 is a solution of the homegenous equation, you do the method of reduction of order by trying to find a solution of the form y= v(x) x^2 and you obtain after plugging it into the equation x^4 v''(x)+x^3 v'(x)=0 x v''(x)+ v'(x)=0 you find v'(x) =K/x so v(x) =ln(x) Hence the second solution is x^2 ln (x) Now the general solution for the homogenous equation is A x^2 + B x^2 ln(x)